a. The ABC company is producing an item at a unit cost of $4. If the supply function is q= 160+8p, where q represents quantity produced and p represents the market price, then


i. Find the total cost function in terms of p


ii. What is the market price at which the company realizes a total profit of $500?



1
Expert's answer
2022-03-18T09:17:29-0400

i. The total cost function in terms of p is:

"TC=Unit Costs\u00d7Q"

"TC=4\u00d7(160+8p)=640+32p"

ii. In order to find market price we have to solve the equation:

"160p+8p^2-640-32p=500"

"8p^2+128p-1140=0"

"D=128\u00d7128-4\u00d78\u00d7(-1140)=52864"

p1= -22.37 (the price can't be negative)

p2=6.37 - the market price at which the company has the profit of $500.